OpenStudy (anonymous):

A random number generator draws at random with replacement from the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. Find the chance that the digit 5 appears on more than 11% of the draws, if:1000 draws are made

OpenStudy (anonymous):

did u get the ans for 100 draws

OpenStudy (anonymous):

the answer is 0.2969668992

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

0.1348v is the ans 4 this ques

OpenStudy (anonymous):

ok I used the graphic calculator and used the binopdf function

OpenStudy (perl):

let X = # fives in 100 random numbers drawn X is a bnomial random variable , p = 1/10 ,

OpenStudy (perl):

it says "more than 11%"

OpenStudy (anonymous):

yes pearl but when I use 1000 I get the answer in calculator is 1 but correct answer is 1.348. Why is it so? will you explain it?

OpenStudy (perl):

wait, what is the answer for the first one?

OpenStudy (perl):

for 1000 draws the answer is .134776

OpenStudy (anonymous):

the answer is 0.2969668992

OpenStudy (anonymous):

yes you are write but when I used the same logic and inset in Ti-84 I got 1.

OpenStudy (perl):

wait,

OpenStudy (perl):

how do you know you got the first part right?

OpenStudy (anonymous):

Great! Pearl I got it. Thanks a lot

OpenStudy (perl):

we want probability picking 5 for more than 11% of the sample (11% of 100 ) so we want P( X > 11 ) P( X > 11 ) = 1 - P(X<=11) = 1 - binomcdf (100, .1, 100*.11)

OpenStudy (perl):

similiarly for 1000 draws we have P( X > 110 ) = 1 - P( X<=110) = 1 - binomcdf ( 1000, .1, 110)

OpenStudy (anonymous):

Yes Bro I got it and understood where I was wrong. Thanks

OpenStudy (perl):

because 11% of 1000 = 110

OpenStudy (perl):

is this online?

OpenStudy (perl):

please tell us if we are right :)

OpenStudy (anonymous):

Yes you are 100% right

OpenStudy (perl):

correct answer is a).2969... b) .1348

OpenStudy (anonymous):

b

OpenStudy (perl):

yeah i was doing the other one from yesterday . any more questions>

OpenStudy (anonymous):

Thanks bro. Not right now